# Giant High-order Nonlinear and Nonreciprocal Electrical Transports Induced by Valley Flipping in Bernal Bilayer Graphene

### v1

@@ Abstract
We investigate the electrical transport properties of the mini-valley polarized state proposed recently in slightly doped Bernal Bilayer Graphene (BLG) in large electric displacement fields. By minimizing the Hartree-Fock energy functional, we first confirm the appearance of mini-valley polarized phase. At the low carrier doping regime, the 1-pocket state will be stabilized where only one of the trigonal-wrapping-induced Fermi pockets near the atomic-valley center is filled. Then we study the electrical transport of the 1-pocket state by solving the Boltzmann equation. We find that the valley polarization could be easily flopped by an in-plane electrical field, which will lead to hysteresis loop in the direct current (DC) $I-V$ curves. Such irreversible current responses in the DC limit will directly induce strong nonlinear and nonreciprocal alternating current (AC) responses, which has been already observed in the recent experiments on BLG.
@@ I Introduction
In the realm of two-dimensional materials, the valley degree of freedom is a critical factor that influences their unique magnetic, transport, and optical characteristics. This is in addition to the well-known degrees of freedom in condensed matter, such as spin, orbital, and momentum. In certain systems, like twisted bilayer graphene and transition metal dichalcogenides (TMDs), a variety of valley polarization and hybridization states have been proposed[1, 2, 3, 4, 5, 6] and observed[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. These states contribute to the emergence of exotic phenomena, including the quantum anomalous Hall effect, correlated insulators, and orbital magnetism. Recently, valley-polarized states have also been detected in non-twisted systems such as Rhombohedral trilayer and Bernal bilayer graphene when a displacement field is applied along the $z$-axis[18, 19, 20, 21, 22]. In this paper, we focus on the Bernal bilayer graphene (BLG) whose atomic structure and Brillouin zone is shown in Fig. 1(a). The displacement field first separates the two Dirac nodes in energy and subsequently leads to the formation of a hybridization gap due to inter-layer coupling[23, 24, 25], resulting in a “Mexican hat” type band dispersion for each atomic valley ±$\pm K$. The trigonal warping effect in BLG further creates three local maxima in the valence band (or minima in the conduction band) near each atomic valley center[24], as indicated by $Q_{1,2,3}$ in Figure 1(b). As a result, the highest valence band (HVB) of gated BLG hosts six band maxima ±$\pm Q_{1,2,3}$ (the prefactor $\xi=\pm$ label the atomic valley, the subscript $\alpha=1,2,3$ label the mini-valley around the atomic valleys), which from now on are called all together as the valleys thought out this paper for simplicity. When the filling factor of the system is just slightly away from the charge neutrality point, quantum oscillation experiments indicate an emergence of the half- and quarter-metal phases[19, 20, 21]. Starting from the quarter-metal phase where only one spin and atomic valley is occupied, if the doping concentration is further reduced so that the mini-valley $Q_{1,2,3}$ could also be viewed as a degree of freedom, a polarized state to the mini-valley is further predicted[26]. Such kind of valley-polarized states break the threefold rotation symmetry $C_{3}$ and has been identified by the $C_{3}$-breaking pattern of angle-resolved second-harmonic nonlinear transport measurements[27]. More recently, anomalous high-order nonlinear and nonreciprocal transport signals have also been detected in the slightly doped BLG[28], challenging traditional transport theories and spurring the need for theoretical investigation from new vantage points.
@@
In this paper, we first establish the presence of mini-valley-polarized phases in gated BLG. This is deduced from the minimization of the system’s total energy functional under Hartree-Fock approximation. The proximity of the three valleys around the $K$ points makes the polarized states quite susceptible to the transitions induced by current flow, potentially leading to significant nonlinear and non-reciprocal transport behaviors. We then analyze the current response of systems exhibiting ”Mexican hat” type band dispersion by employing the Boltzmann equation, incorporating the effects of long-range Coulomb interactions among carriers. Differentiating between relaxation processes within the same mini-valley and across different valleys allowed us to derive equations that effectively describe the dynamics within these systems. The resulting phenomena from our analysis include: 1) The ease of flipping mini-valley polarization via an electric current leading to hysteresis loops in the direct current (DC) $I-V$ profile. 2) The presence of an irreversible current response in the DC regime leads to significant nonlinear and nonreciprocal alternating current (AC) responses at low frequencies, where the current response remains adiabatic to the external drive. 3) The existence of a critical external field intensity where the long-term behavior of the current response becomes largely insensitive to the initial polarization direction, leading to a near-symmetrical angular response in the system.
The rest of the paper will be organized as follows. In Section II, we first confirm the existence of the mini-valley polarized ground states in the very low hole doping regime by minimizing the Hartree Fock mean-field energy functional. In Section III, we solve the Boltzmann equation for the mini-valley polarized ground state with a DC electrical field, and a valley flipping phenomenon is identified when the electrical field strength reaches some critical value. By replacing the DC electrical field by a low frequency sinusoidal AC field, the AC electrical transports are investigated in Section IV. We find that the system enters into a steady valley oscillating state if the AC electrical field amplitude is larger that the DC critical field for valley flipping. At this time, the current responses in time domain show discontinuous features which lead to very high order nonlinearity in frequency domain responses. Besides, we also find that the valley oscillating state is insensitive to the initial polarization direction when the electrical field amplitude is large enough, and the current responses show a $C_{3}$-persevering angular pattern although the initial mini-valley polarized ground state breaks the threefold rotation symmetry. In Section V, symmetry restrictions on the angular pattern of the current responses are discussed in more details.
@@ II The Valley-Polarized Ground States
Under Hartree-Fock approximation, the total energy per area is expressed as a functional of the hole occupation function $n_{\bm{k}}\equiv 1-\langle c^{\dagger}_{\bm{k}}c_{\bm{k}}\rangle$ as $E_{tot}[n_{\bm{k}}]=E_{0}+E_{H}+E_{K}[n_{\bm{k}}]+E_{ex}[n_{\bm{k}}]$. The vacuum energy of the fully filled HVB $E_{0}=\mathcal{S}^{-1}\sum_{\bm{k}}\varepsilon^{0}_{\bm{k}}$ is a constant. Besides, we assume the system is charge uniform and the Hartree energy $E_{H}$ is also a constant. Thus these two terms will be ignored in the following text. The other two terms are kinetic energy $E_{K}$ and exchange energy $E_{ex}$, which are separately written as
@@
Then we can treat the three $n_{\alpha}$s as the only variational parameters to greatly simplify the calculations. In Fig. 2(b-d), the approximated hole occupation function gotten by Eq. (4) (shaded areas) are compared with those from the full variational scheme (red lines), which reveal a significant degree of agreement.
@@
$E_{tot}[n_{\bm{k}}]\approx E_{tot}(n_{1},n_{2},n_{3}).$ (6)
More details of the energy functional is given in Sec. I of the supplemental materials (SM). By minimizing Eq. (6) with respect to $n_{\alpha}$ under constrain $n_{1}+n_{2}+n_{3}=n_{h}$, we find that the only possible stable sates (local minimums) are the $N$-pocket states ($N=1,2,3$), where there are $N$ hole pockets equally occupied. And this is consistent with the previous study by Dong et al. [26]. In Fig. 2(a), the total energies per hole $E_{tot}/n_{h}$ of the $N$-pocket states are plotted in the hole density range where they are stable, and the ground state is the one with the lowest energy. The typical ground state hole distributions of the $N$-pocket states are also plotted in Fig. 2(b-d).
@@ III DC Transport: Valley Flipping and Current Hysteresis
Figure 3: (a) Starting from the $\mathcal{Q}_{1}$-polarized state $P_{0}$ at zero electrical field strength, the steady hole distributions are smoothly connected if the electrical fields vary in the blue region and its boundary indicates the critical electrical field for valley flipping. (b)(c) The same quantities as (a) except that the initial states at zero field strength are polarized in the $\mathcal{Q}_{2}$ and $\mathcal{Q}_{3}$ valleys respectively. (d-f) Steady hole distributions for points $P_{0-4}$ in (a). (g-i) Steady hole distributions for points $P_{0-4}^{\prime}$ in (b). The blue and red color boxed panels in (d-i) form a hysteresis loop when the electrical field varies along the blue and red lines in (a) and (b). When generating these plots, the parameters are taken as $n_{h}=0.06\times 10^{12}\;\mathrm{cm}^{-2}$, $\tau_{i}=10\;\mathrm{ps}$ and $\tau_{x}/\tau_{i}=100$. The field unit is defined as $E^{*}=\hbar\sqrt{3}Q/(2e\tau_{i})\approx 19.4\mathrm{mV}/\mathrm{\mu m}$. The direction of the electrical field indicated by the blue and red lines is $\theta=7\pi/18$.
@@
The first term on the right side of the equation describes the driving effect by a DC electrical field and the second term is the impurity scattering, where $n_{imp}$ is the impurity concentration, $T_{\bm{k}\bm{k}^{\prime}}$ is the impurity scattering matrix and $\tilde{\varepsilon}_{\bm{k}}\equiv\partial_{n_{\bm{k}}}E_{tot}[n_{\bm{k}}]$ is the quasi-particle energy of the holes. Considering Eq. (3)(4), the quai-particle energy in the $\mathcal{Q}_{\alpha}$ valley could be approximated as $\tilde{\varepsilon}_{\bm{k}}\approx-\varepsilon^{0}_{\bm{k}}+\Sigma_{\bm{k}}$ where $\Sigma_{\bm{k}}\approx-{\mathcal{S}}^{-1}\sum_{\bm{k}^{\prime}}U_{\bm{k}\bm{k}^{\prime}}(\bm{k}-\bm{k}^{\prime})n^{0}_{\bm{k}^{\prime}}$ is the self-energy of holes due to interaction. Near the mini-valley center, the self-energy could be viewed as constants $\Sigma_{\bm{k}}\approx\Sigma_{\alpha[\bm{k}]}$ for different valleys, where $\Sigma_{\alpha}$ is obtained by $\Sigma_{\alpha}\approx\partial_{n_{\alpha}}E_{ex}(n_{1},n_{2},n_{3})$. As for the scattering matrix, we also assume that $|T_{\bm{k}\bm{k}^{\prime}}|$ to be constants $T_{i}$ and $T_{x}$ for the intra- and inter-valley scattering process, which satisfies $T_{i}\gg T_{x}$. For charge impurity with long range impurity potential in real space, the scattering matrix drops quickly with respect to the transferred momentum $|\bm{k}-\bm{k}^{\prime}|$, and $T_{i}\gg T_{x}$ wil be a good approximation. Under the above assumption, the intra-valley relaxation process is considered to be much faster than the inter-valley process. Therefore, we can approximately separate the fast and slow processes by ignoring the inter-valley relaxation process when dealing with the intra-valley dynamics. Assume the hole concentration in $\mathcal{Q}_{\alpha}$ valley is $n_{\alpha}$ at time $t$, by adopting the relaxation time approximation, the Boltzmann equation in $\mathcal{Q}_{\alpha}$ valley becomes
@@
Notice that $\dot{\bm{k}}=e\bm{E}/\hbar$ is the velocity of in $k$-space, we can define the current flows in $k$-space as $\bm{\mathcal{J}}_{\bm{k}}\equiv\dot{\bm{k}}n_{\bm{k}}$. Using Gauss’ law, Eq. (11) could be further written as
@@
In other words, the field driving term could be interpreted as changing rate of the particle numbers in area $\mathcal{Q}_{\alpha}$ caused by the $k$-space current flow $\bm{\mathcal{J}}_{\bm{k}}$. The second term on the right side of Eq. (10) is the changing rate in area $\mathcal{Q}_{\alpha}$ due to impurity scattering. According to the derivation in Sec. II of the SM, this term could be approximated by
@@
In Fig. 3 and 4, the hole density is taken as $n_{h}=0.06\times 10^{12}\;\mathrm{cm}^{-2}$ and the relaxation times are taken as $\tau_{i}=10\;\mathrm{ps}$ and $\tau_{x}/\tau_{i}=100$. The electrical field unit is defined as $E^{*}=\hbar\sqrt{3}Q/(2e\tau_{i})\approx 19.4\mathrm{mV}/\mathrm{\mu m}$, where $Q=|\bm{Q}_{\alpha}-\bm{K}_{+}|$. The current unit is defined by $j^{*}=e^{2}\tau_{i}n_{h}E^{*}/\sqrt{m_{r}m_{\varphi}}\approx 3.33\;\mathrm{nA}/\mathrm{nm}$ where $m_{r}$ and $m_{\varphi}$ are effective masses at the mini-valley center in radial and angular direction respectively.

### v2

@@ Abstract
We investigate the electrical transport properties of the mini-valley polarized state proposed recently in slightly doped Bernal Bilayer Graphene (BLG) in large electric displacement fields. By minimizing the Hartree-Fock (HF) energy functional, we first confirm the appearance of mini-valley polarized phase. At the low carrier doping regime, the 1-pocket state will be stabilized where only one of the trigonal-wrapping-induced Fermi pockets near the atomic-valley center is filled. Then we study the electrical transport of the 1-pocket state by solving the Boltzmann equation. We find that the valley polarization could be easily flopped by an in-plane electrical field, which will lead to hysteresis loop in the direct current (DC) I-V curves. Such irreversible current responses in the DC limit will directly induce strong nonlinear and nonreciprocal alternating current (AC) responses, which has been already observed in the recent experiments on BLG.
pacs:
@@ I Introduction
In the realm of two-dimensional materials, the valley degree of freedom is a critical factor that influences their unique magnetic, transport, and optical characteristics. This is in addition to the well-known degrees of freedom in condensed matter, such as spin, orbital, and momentum. In certain systems, like twisted bilayer graphene (TBG) and transition metal dichalcogenides (TMDs), a variety of valley polarization and hybridization states have been proposed[1, 2, 3, 4, 5, 6] and observed[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. These states contribute to the emergence of exotic phenomena, including the quantum anomalous Hall effect, correlated insulators, and orbital magnetism. Recently, valley-polarized states have also been detected in non-twisted systems such as Rhombohedral trilayer and Bernal bilayer graphene when a displacement field is applied along the $z$-axis[18, 19, 20, 21, 22]. In this paper, we focus on the Bernal bilayer graphene (BLG) whose atomic structure and Brillouin zone is shown in Fig. 1(a). The displacement field first separates the two Dirac nodes in energy and subsequently leads to the formation of a hybridization gap due to inter-layer coupling[23, 24, 25], resulting in a “Mexican hat” type band dispersion for each atomic valley ±$\pm K$. The trigonal warping effect in BLG further creates three local maxima in the valence band (or minima in the conduction band) near each atomic valley center[24], as indicated by $Q_{1,2,3}$ in Figure 1(b). As a result, the highest valence band (HVB) of gated BLG hosts six band maxima ±$\pm Q_{1,2,3}$ (the prefactor $\xi=\pm$ label the atomic valley, the subscript $\alpha=1,2,3$ label the mini-valley around the atomic valleys), which from now on are called all together as the valleys thought out this paper for simplicity. When the filling factor of the system is just slightly away from the charge neutrality point, quantum oscillation experiments indicate an emergence of the half- and quarter-metal phases[19, 20, 21]. Starting from the quarter-metal phase where only one spin and atomic valley is occupied, if the doping concentration is further reduced so that the mini-valley $Q_{1,2,3}$ could also be viewed as a degree of freedom, a polarized state to the mini-valley is further predicted[26]. Such kind of valley-polarized states break the threefold rotation symmetry $C_{3}$ and has been identified by the $C_{3}$-breaking pattern of angle-resolved second-harmonic nonlinear transport measurements[27]. More recently, anomalous high-order nonlinear and nonreciprocal transport signals have also been detected in the slightly doped BLG[28], challenging traditional transport theories and spurring the need for theoretical investigation from new vantage points.
@@
In this paper, we first establish the presence of mini-valley-polarized phases in gated BLG. This is deduced from the minimization of the system’s total energy functional under Hartree-Fock (HF) approximation. The proximity of the three valleys around the $K$ points makes the polarized states quite susceptible to the transitions induced by current flow, potentially leading to significant nonlinear and non-reciprocal transport behaviors. We then analyze the current response of systems exhibiting ”Mexican hat” type band dispersion by employing the Boltzmann equation, incorporating the effects of long-range Coulomb interactions among carriers. Differentiating between relaxation processes within the same mini-valley and across different valleys allowed us to derive equations that effectively describe the dynamics within these systems. The resulting phenomena from our analysis include: (1) The mini-valley polarization can be reversed via electric current, causing first-order hysteresis loops in the direct current (DC) I-V curve, which was also pointed out in a recent study[42]. (2) The emergence of irreversible current responses in the DC spectrum, resulting in significant nonlinear and nonreciprocal alternating current (AC) behaviors at low frequencies. (3) The presence of a critical field intensity threshold beyond which the system’s long-term current response becomes insensitive to initial polarization directions, yielding nearly symmetric angular responses.
We stress that the first-order hysteretic behaviors in the DC I-V curves and the related phenomena discussed above persist as long as the ground state is mini-valley polarized. This means that these phenomena are also “smoking gun” for confirming the mini-valley polarized ground state.
Similar current-induced hysteretic phenomena were also observed[12, 8] and studied[29, 30] in TBG. Nonetheless, the mechanisms in TBG are quite different from this work. Firstly, the ground state in TBG is an insulator while the ground state in this study is a metal. Besides, as will be shown in Sec.III, the valley flipping phenomenon in our work results from the competition between the electric field driving effect and the inter-valley relaxation. However, neither of these two processes could possibly exist in any of the insulator phase, such as the previous theoretical studies in TBG[29, 30]. In other words, we present a new mechanism of current-induced valley polarization flipping phenomenon in the valley polarized metal.
@@ II The Valley-Polarized Ground States
Under HF approximation, the total energy per area is expressed as a functional of the hole occupation function $n_{\bm{k}}\equiv 1-\langle c^{\dagger}_{\bm{k}}c_{\bm{k}}\rangle$ as $E_{tot}[n_{\bm{k}}]=E_{0}+E_{H}+E_{K}[n_{\bm{k}}]+E_{ex}[n_{\bm{k}}]$. The vacuum energy of the fully filled HVB $E_{0}=\mathcal{S}^{-1}\sum_{\bm{k}}\varepsilon^{0}_{\bm{k}}$ is a constant. Besides, we assume the system is charge uniform and the Hartree energy $E_{H}$ is also a constant. Thus these two terms will be ignored in the following text. The other two terms are kinetic energy $E_{K}$ and exchange energy $E_{ex}$, which are separately written as
@@
Then we can treat the three $n_{\alpha}$s as the only variational parameters to greatly simplify the calculations. In Fig. 2(b-d), the approximated hole occupation function gotten by Eq. (4) (shaded areas) are compared with those from the full variational scheme (red lines), which reveal a significant degree of agreement. Besides, the total energies calculated from the approximated occupation function Eq. (4) are also compared with the results from the full variational scheme in Appendix A, and the relative errors are in the order of $10^{-3}$. These two facts indicate that Eq. (4) is a very good approximation of the occupation function.
@@
$E_{tot}[n_{\bm{k}}]\approx E_{K}(n_{1},n_{2},n_{3})+E_{ex}(n_{1},n_{2},n_{3}).$ (6)
For numerically convenience, the exchange energy functional are separated into the intra- and inter-valley terms which are further approximated by Eq. (21) and Eq. (22) respectively. By minimizing Eq. (6) with respect to $n_{\alpha}$ under constrain $n_{1}+n_{2}+n_{3}=n_{h}$, we find that the only possible stable sates (local minimums) are the $N$-pocket states ($N=1,2,3$), where there are $N$ hole pockets equally occupied. And this is consistent with the previous study by Dong et al. [26]. In Fig. 2(a), the total energies per hole $E_{tot}/n_{h}$ of the $N$-pocket states are plotted in the hole density range where they are stable, and the ground state is the one with the lowest energy. The typical ground state hole distributions of the $N$-pocket states are also plotted in Fig. 2(b-d).
Figure 3: The critical density between the 1- and 2-pocket states $n_{c,12}$ and the critical density between the 2- and 3-pocket states $n_{c,23}$.
Denote the critical density between the 1- and 2-pocket states as $n_{c,12}$, the critical density between the 2- and 3-pocket states as $n_{c,23}$. In Fig. 3, the critical densities are plotted as functions of the inter-layer bias potential $U$. The corresponding displacement field $D=U/d$ ($d$ is the inter-layer distance) is also marked in the right axis for reference. At $U=200\;\mathrm{meV}$ and $D\approx 600\;\mathrm{mV}/\mathrm{nm}$, the critical density for the 1- and 2-pocket state is about $n_{c,12}\approx 0.11\times 10^{12}\;\mathrm{cm}^{-2}$, which is a little larger than the experimentally result $0.07\times 10^{12}\;\mathrm{cm}^{-2}$ measured by Lin et al. [27]. The overestimation is reasonable since the screening effect of the interaction is not included in the Hartree Fock approximation.
@@ III DC Transport: Valley Flipping and Current Hysteresis
Figure 4: (a) Starting from the $\mathcal{Q}_{1}$-polarized state $P_{0}$ at zero electrical field strength, the steady hole distributions are smoothly connected if the electrical fields vary in the blue region and its boundary indicates the critical electrical field for valley flipping. (b)(c) The same quantities as (a) except that the initial states at zero field strength are polarized in the $\mathcal{Q}_{2}$ and $\mathcal{Q}_{3}$ valleys respectively. (d-f) Steady hole distributions for points $P_{0-4}$ in (a). (g-i) Steady hole distributions for points $P_{0-4}^{\prime}$ in (b). The blue and red color boxed panels in (d-i) form a hysteresis loop when the electrical field varies along the blue and red lines in (a) and (b). When generating these plots, the parameters are taken as $n_{h}=0.06\times 10^{12}\;\mathrm{cm}^{-2}$, $\tau_{i}=100\;\mathrm{fs}$ and $\tau_{x}/\tau_{i}=100$. The field unit is defined as $E^{*}=\hbar\sqrt{3}Q/(2e\tau_{i})\approx 1.94\;\mathrm{V}/\mathrm{\mu m}$. The direction of the electrical field indicated by the blue and red lines is $\theta=7\pi/18$.
@@
The first term on the right side of the equation describes the driving effect by a DC electrical field and the second term is the impurity scattering, where $n_{imp}$ is the impurity concentration, $T_{\bm{k}\bm{k}^{\prime}}$ is the impurity scattering matrix and $\tilde{\varepsilon}_{\bm{k}}\equiv\partial_{n_{\bm{k}}}E_{tot}[n_{\bm{k}}]$ is the quasi-particle energy of the holes. Considering Eq. (3)(4), the quai-particle energy in the $\mathcal{Q}_{\alpha}$ valley could be approximated as $\tilde{\varepsilon}_{\bm{k}}\approx-\varepsilon^{0}_{\bm{k}}+\Sigma_{\bm{k}}$ where $\Sigma_{\bm{k}}\approx-{\mathcal{S}}^{-1}\sum_{\bm{k}^{\prime}}U_{\bm{k}\bm{k}^{\prime}}(\bm{k}-\bm{k}^{\prime})n^{0}_{\bm{k}^{\prime}}$ is the self-energy of holes due to interaction (More precisely, the self-energy should be $\Sigma_{\bm{k}}=-{\mathcal{S}}^{-1}\sum_{\bm{k}^{\prime}}U_{\bm{k}\bm{k}^{\prime}}(\bm{k}-\bm{k}^{\prime})n_{\bm{k}^{\prime}}$, and the solutions of the Boltzmann equation with/without this approximation are compared in appendix. D, which only have little difference). Near the mini-valley center, the self-energy could be viewed as constants $\Sigma_{\bm{k}}\approx\Sigma_{\alpha[\bm{k}]}$ for different valleys, where $\Sigma_{\alpha}$ is obtained by $\Sigma_{\alpha}\approx\partial_{n_{\alpha}}E_{ex}(n_{1},n_{2},n_{3})$. As for the scattering matrix, we also assume that $|T_{\bm{k}\bm{k}^{\prime}}|$ to be constants $T_{i}$ and $T_{x}$ for the intra- and inter-valley scattering process, which satisfies $T_{i}\gg T_{x}$. For charge impurity with long range impurity potential in real space, the scattering matrix drops quickly with respect to the transferred momentum $|\bm{k}-\bm{k}^{\prime}|$, and $T_{i}\gg T_{x}$ wil be a good approximation. Under the above assumption, the intra-valley relaxation process is considered to be much faster than the inter-valley process. Therefore, we can approximately separate the fast and slow processes by ignoring the inter-valley relaxation process when dealing with the intra-valley dynamics. Assume the hole concentration in $\mathcal{Q}_{\alpha}$ valley is $n_{\alpha}$ at time $t$, by adopting the relaxation time approximation, the Boltzmann equation in $\mathcal{Q}_{\alpha}$ valley becomes
@@
Notice that $\dot{\bm{k}}=e\bm{E}/\hbar$ is the velocity in $k$-space, we can define the current flows in $k$-space as $\bm{\mathcal{J}}_{\bm{k}}\equiv\dot{\bm{k}}n_{\bm{k}}$. Using Gauss’ law, Eq. (11) could be further written as
@@
In other words, the field driving term could be interpreted as changing rate of the particle numbers in area $\mathcal{Q}_{\alpha}$ caused by the $k$-space current flow $\bm{\mathcal{J}}_{\bm{k}}$. The second term on the right side of Eq. (10) is the changing rate in area $\mathcal{Q}_{\alpha}$ due to impurity scattering. According to the derivation in Appendix C, this term could be approximated by
@@
In Fig. 4 and 5, the hole density is taken as $n_{h}=0.06\times 10^{12}\;\mathrm{cm}^{-2}$ and the relaxation times are taken as $\tau_{i}=100\;\mathrm{fs}$[40] and $\tau_{x}/\tau_{i}=100$. The electrical field unit is defined as $E^{*}=\hbar\sqrt{3}Q/(2e\tau_{i})\approx 1.94\;\mathrm{V}/\mathrm{\mu m}$, where $Q=|\bm{Q}_{\alpha}-\bm{K}_{+}|$. The current unit is defined by $j^{*}=e^{2}\tau_{i}n_{h}E^{*}/\sqrt{m_{r}m_{\varphi}}\approx 3.33\;\mathrm{nA}/\mathrm{nm}$ where $m_{r}$ and $m_{\varphi}$ are effective masses at the mini-valley center in radial and angular direction respectively.